# SRIF: Semantic Shape Registration Empowered by Diffusion-based Image Morphing and Flow Estimation

### v1

@@ 4.1. Implementation Details
For the input source and target shape, we use Open3D to sample uniformly $16$ viewpoints. In particular, empirically we render the background of all rendered images to black for the best overall performance. Note that we demonstrate a white background in Fig. 3, which is for better visualization. We use DiffMorpher [342] to interpolate $10$ frames between each pair of images rendered from the same viewpoint. Regarding SC-GS [130], we use vertices extracted from the source mesh as the initialization of the mean of Gaussians. The first $3000$ iterations of the training step are used for initializing each single Gaussian. Then we train a model to predict the deformation between them, simultaneously optimizing the position, opacity, and covariance matrix for a total of $20000$ iterations. We set the percent dense parameter to $0.01$ to generate relatively sparse point clouds. In general, we believe in target shape more than the intermediate point clouds. During flow estimation, we set the weight $\lambda_{\mbox{cd\_inter}}$ to $1$, that of $\lambda_{\mbox{cd\_final}}$ to $10$, and $\lambda_{\mbox{arap}}$ to $2$ for all categories. We train $4000$ iterations per pair and the learning rate is set to $31e-$ 3. We set $t_{0}=0$, $t_{1}=0.5$ and using the $2nd,4th,6th$ and $8th$ of the $T=10$ reconstructed point clouds in Eqn. 3.
Table 1. Average geodesic errors on each of the test categories in SHREC07 [103] and EBCM [328].
@@ 4.2. Evaluation Setup
Baselines: In this section, we compare our method with an array of methods of estimating dense correspondences, which 1) require no external landmarks as input and 2) pose no constraint (e.g., near-isometry) on the deformations between input shapes: BIM [144], SmoothShells [83], NeuroMorph [84], MapTree [242], and ENIGMA [82]
Benchmarks: We comprehensively compare our method and the baselines in an array of test sets. (1) We consider $9$ categories from SHREC07 dataset [103]human, fourleg, airplane, bird, chair, fish, ant, pier, and glasses, each of which contains $20$ shapes. In order to fully evaluate the capacity of all methods, we select the most distinctive shape pairs via the following scheme. For a pair of shapes $S_{i},S_{j}$ from the same category, we compute the first $50$ eigenvalues $\Lambda_{i},\Lambda_{j}$ and the spectral distance $d(S_{i},S_{j})=\|\Lambda_{i}-\Lambda_{j}\|$. We then sample $50$ pairs from the overall $190$ in the descending order of spectral distances for our test. (2) We further pick ten pairs from the test cases presented in [328] for more variability. Finally, we remark that for each pair $(S_{i},S_{j})$, we compute maps in both directions.
Evaluation Metrics: To evaluate the maps, we consider four well-known metrics including Dirichlet energy [90], Coverage (i.e. surjectivity) [125], Geodesic matching error with respect to ground truth landmark labels [144], and Bijectivity [245]. We refer readers to the Supp. Mat. for the details on these metrics.
@@ 4.4. Dense Shape Correspondence
In total, SRIF achieves an average Geodesic Error of $0.0956$, which is less than half of the second-best baseline. Our method even outperforms ENIGMA in Dirichlet Energy – our method does not explicitly optimize for this metric while ENIGMA uses RHM [90] for post-processing, which minimizes Dirichlet energy. Fig. 10 showcases part of the qualitative results, which agree nicely with the quantitative results. One obvious problem among the intrinsic-based method (BIM, MapTree, Enigma) is symmetry ambiguity. Though our method gets rid of such by extrinsic alignment, we argue that such a requirement is indeed easier to meet than injecting orientation information into intrinsic-based methods. Moreover, in the Supp. Mat., we also report the scores of ENIGMA with allowance on the symmetric flip (as in the regarding paper), yet our method still outperforms it across all sets with more restrictive evaluation.
@@ 4.6. Robustness
Mesh Quality: In practice, meshes can exhibit severe irregularities. For instance, the shapes in the right-most and left-most columns of Fig. 8 are from KeyPointNet dataset [333], which consist of a large portion of thin triangles. This can pose significant challenges on registration methods based on triangulation and point-wise correspondence update. For instance, in the top row of Fig. 8, SmoothShells cannot match the airplanes even though they are similar to each other. In contrast, thanks to the flow estimation module, our method enjoys stronger robustness on this front.
@@
Rotational Perturbation: We generally assume the shapes of interest are extrinsically aligned, which is a common practice in shape registration [331]. Meanwhile, we empirically observe that SRIF is robust with rotational perturbation as large as $45$ degrees (see Tab. 3 and qualitative results in the Supp. Mat.). We attribute such to the fact that the diffusion model is trained with objects in various orientations [86], thus gaining certain robustness and passing to our framework.
@@ 4.7. Ablation Studies
Reconstruction Method: We perform dynamic 3D-GS on the whole set of multi-view interpolated images. One can as well independently perform 3D Gaussian Splatting [138] on the multi-view images at each time step to obtain the intermediate geometry. The third row of Tab. 4 suggests that the above variant is suboptimal. This is probably since image interpolation is performed independently from each viewpoint, therefore it is hard to guarantee the multi-view consistency at each time step.
@@ 4.8. Running Time Analysis
Figure 10. Qualitative comparisons on various categories from SHREC’07 [103]. ENIGMA [82] fails to return result on the pair of birds. Note that the target chair in the second row is not disconnected – It appears so due to the thin structure within the chair.
@@ 5. Conclusion and Limitations
We identify the following limitations of our method, which lead to future work directions: 1) There exists significant room for improvement on efficiency. As shown in Sec. 4.8, the main bottleneck of our pipeline lies in image morphing, or, more specifically, LoRA fine-tuning, which takes over $50\%$ of the total running time. To this end, we plan to resort to advances in parameter efficient fine-tuning; 2) Our method does not guarantee the output to be continuous or bijective. It would be interesting to explore how to regularize the smoothness of flow [78]; on the other hand, since flow is by construction invertible, we can encourage bijectivity by taking forward and backward flow simultaneously during training; 3) Since our method leverages image interpolation, it could be suboptimal when the intermediate results are problematic. To see that, we evaluate our method on SHREC19 [191] and SMPL [178] dataset. As shown in Tab. 5, our method is outperformed by SmoothShells, which leverages intrinsic geometry information, with a notable margin. We attribute the failure to two factors: first human articulation spans a large space, which might not be well learnt by diffusion models without explicit modeling; second there exist self-occlusions among the multi-view renderings of human shapes. We refer readers to the Supp. Mat. for more details on this experiment. We plan to introduce stronger prior on this front, say, leveraging pre-trained model on human shapes/images. 4) Our method uses ARAP to regularize mesh deformation, which implicitly encourages local neighborhood preservation. This in turn prevents our method from characterizing significant topological changes (see bottom row of Fig. 8, where the supports of tables are distinctive). As shown in Sec. 4.5, our method can be adapted into purely point-based settings, it is interesting to further exploit this property.
@@ A.6. Non-rigid shape matching
In this section, we discuss the performance of our method on datasets with significant pose variations. We selected the following two datasets: For SHREC19 [191], we evaluate 407 pairs of data with ground truth. In particular, we exclude 23 pairs related to a partial shape. For SMPL [178], we randomly generate a set of 500 shapes using SMPL model. And then sample 20 shapes from them via FPS in the pose parameters of generation. Subsequently, similar to the procedure in Sec.4.2 of the main text, we construct 50 pairs among the 20 shapes for inference. As shown in Tab. 5 of the main text, our method is outperformed by SmoothShells with a noticeable margin.
To investigate the failure cause, we examine the intermediate results of our pipeline. As shown in Fig. 15, when there are significant differences in the pose, the image interpolation module [342] struggles to return plausible results (top row), which is further amplified in the follow-up point cloud reconstruction (bottom row). Such discrepancy then leads to the suboptimal solution of our method on this benchmark.
@@ A.7. Running Time analysis
In particular, we evaluate MapTree  [^1] https://github.com/llorz/SGA20_mapExplor, SmoothShells  [^2] https://github.com/marvin-eisenberger/smooth-shells, NeuroMorph  [^3] https://github.com/facebookresearch/neuromorph, and BIM on all the test data with the regarding official implementations. Regarding ENIGMA, since the code is not publicly available, we asked the authors to run baseline evaluation, who also reported that ENIGMA took 20 mins to match shapes of 5000 vertices with RHM [90] as post-refinement.

### v2

@@ 4.1. Implementation Details
For the input source and target shape, we use Open3D to sample uniformly $16$ viewpoints. In particular, empirically we render the background of all rendered images to black for the best overall performance. Note that we demonstrate a white background in Fig. 3, which is for better visualization. We use DiffMorpher (Zhang et al., 2023) to interpolate $10$ frames between each pair of images rendered from the same viewpoint. Regarding SC-GS (Huang et al., 2024), we use vertices extracted from the source mesh as the initialization of the mean of Gaussians. The first $3000$ iterations of the training step are used for initializing each single Gaussian. Then we train a model to predict the deformation between them, simultaneously optimizing the position, opacity, and covariance matrix for a total of $20000$ iterations. We set the percent dense parameter to $0.01$ to generate relatively sparse point clouds. In general, we believe in target shape more than the intermediate point clouds. During flow estimation, we set the weight $\lambda_{\mbox{cd\_inter}}$ to $1$, that of $\lambda_{\mbox{cd\_final}}$ to $10$, and $\lambda_{\mbox{arap}}$ to $2$ for all categories. We train $4000$ iterations per pair and the learning rate is set to $31e-$ 3. We set $t_{0}=0$, $t_{1}=0.5$ and using the $2nd,4th,6th$ and $8th$ of the $T=10$ reconstructed point clouds in Eqn. 3.
Table 1. Average geodesic errors on each of the test categories in SHREC07 (Giorgi et al., 2007) and EBCM (Yang et al., 2020).
@@ 4.2. Evaluation Setup
Baselines: In this section, we compare our method with an array of methods of estimating dense correspondences, which 1) require no external landmarks as input and 2) pose no constraint (e.g., near-isometry) on the deformations between input shapes: BIM (Kim et al., 2011), SmoothShells (Eisenberger et al., 2020), NeuroMorph (Eisenberger et al., 2021), MapTree (Ren et al., 2020), and ENIGMA (Edelstein et al., 2020)
Benchmarks: We comprehensively compare our method and the baselines in an array of test sets. (1) We consider $9$ categories from SHREC07 dataset (Giorgi et al., 2007)human, fourleg, airplane, bird, chair, fish, ant, pier, and glasses, each of which contains $20$ shapes. In order to fully evaluate the capacity of all methods, we select the most distinctive shape pairs via the following scheme. For a pair of shapes $S_{i},S_{j}$ from the same category, we compute the first $50$ eigenvalues $\Lambda_{i},\Lambda_{j}$ and the spectral distance $d(S_{i},S_{j})=\|\Lambda_{i}-\Lambda_{j}\|$. We then sample $50$ pairs from the overall $190$ in the descending order of spectral distances for our test. (2) We further pick ten pairs from the test cases presented in (Yang et al., 2020) for more variability. Finally, we remark that for each pair $(S_{i},S_{j})$, we compute maps in both directions.
Evaluation Metrics: To evaluate the maps, we consider four well-known metrics including Dirichlet energy (Ezuz et al., 2019), Coverage (i.e. surjectivity) (Huang and Ovsjanikov, 2017), Geodesic matching error with respect to ground truth landmark labels (Kim et al., 2011), and Bijectivity (Ren et al., 2018). We refer readers to the Supp. Mat. for the details on these metrics.
@@ 4.4. Dense Shape Correspondence
In total, SRIF achieves an average Geodesic Error of $0.0956$, which is less than half of the second-best baseline. Our method even outperforms ENIGMA in Dirichlet Energy – our method does not explicitly optimize for this metric while ENIGMA uses RHM (Ezuz et al., 2019) for post-processing, which minimizes Dirichlet energy. Fig. 10 showcases part of the qualitative results, which agree nicely with the quantitative results. One obvious problem among the intrinsic-based method (BIM, MapTree, Enigma) is symmetry ambiguity. Though our method gets rid of such by extrinsic alignment, we argue that such a requirement is indeed easier to meet than injecting orientation information into intrinsic-based methods. Moreover, in the Supp. Mat., we also report the scores of ENIGMA with allowance on the symmetric flip (as in the regarding paper), yet our method still outperforms it across all sets with more restrictive evaluation.
@@ 4.6. Robustness
Mesh Quality: In practice, meshes can exhibit severe irregularities. For instance, the shapes in the right-most and left-most columns of Fig. 8 are from KeyPointNet dataset (You et al., 2020), which consist of a large portion of thin triangles. This can pose significant challenges on registration methods based on triangulation and point-wise correspondence update. For instance, in the top row of Fig. 8, SmoothShells cannot match the airplanes even though they are similar to each other. In contrast, thanks to the flow estimation module, our method enjoys stronger robustness on this front.
@@
Rotational Perturbation: We generally assume the shapes of interest are extrinsically aligned, which is a common practice in shape registration (Yao et al., 2023). Meanwhile, we empirically observe that SRIF is robust with rotational perturbation as large as $45$ degrees (see Tab. 3 and qualitative results in the Supp. Mat.). We attribute such to the fact that the diffusion model is trained with objects in various orientations (El Banani et al., 2024), thus gaining certain robustness and passing to our framework.
@@ 4.7. Ablation Studies
Reconstruction Method: We perform dynamic 3D-GS on the whole set of multi-view interpolated images. One can as well independently perform 3D Gaussian Splatting (Kerbl et al., 2023) on the multi-view images at each time step to obtain the intermediate geometry. The third row of Tab. 4 suggests that the above variant is suboptimal. This is probably since image interpolation is performed independently from each viewpoint, therefore it is hard to guarantee the multi-view consistency at each time step.
@@ 4.8. Running Time Analysis
Figure 10. Qualitative comparisons on various categories from SHREC’07 (Giorgi et al., 2007). ENIGMA (Edelstein et al., 2020) fails to return result on the pair of birds. Note that the target chair in the second row is not disconnected – It appears so due to the thin structure within the chair.
@@ 5. Conclusion and Limitations
We identify the following limitations of our method, which lead to future work directions: 1) There exists significant room for improvement on efficiency. As shown in Sec. 4.8, the main bottleneck of our pipeline lies in image morphing, or, more specifically, LoRA fine-tuning, which takes over $50\%$ of the total running time. To this end, we plan to resort to advances in parameter efficient fine-tuning; 2) Our method does not guarantee the output to be continuous or bijective. It would be interesting to explore how to regularize the smoothness of flow (Dupuis et al., 1998); on the other hand, since flow is by construction invertible, we can encourage bijectivity by taking forward and backward flow simultaneously during training; 3) Since our method leverages image interpolation, it could be suboptimal when the intermediate results are problematic. To see that, we evaluate our method on SHREC19 (Melzi et al., 2019) and SMPL (Loper et al., 2015) dataset. As shown in Tab. 5, our method is outperformed by SmoothShells, which leverages intrinsic geometry information, with a notable margin. We attribute the failure to two factors: first human articulation spans a large space, which might not be well learnt by diffusion models without explicit modeling; second there exist self-occlusions among the multi-view renderings of human shapes. We refer readers to the Supp. Mat. for more details on this experiment. We plan to introduce stronger prior on this front, say, leveraging pre-trained model on human shapes/images. 4) Our method uses ARAP to regularize mesh deformation, which implicitly encourages local neighborhood preservation. This in turn prevents our method from characterizing significant topological changes (see bottom row of Fig. 8, where the supports of tables are distinctive). As shown in Sec. 4.5, our method can be adapted into purely point-based settings, it is interesting to further exploit this property.
@@ A.6. Non-rigid shape matching
In this section, we discuss the performance of our method on datasets with significant pose variations. We selected the following two datasets: For SHREC19 (Melzi et al., 2019), we evaluate 407 pairs of data with ground truth. In particular, we exclude 23 pairs related to a partial shape. For SMPL (Loper et al., 2015), we randomly generate a set of 500 shapes using SMPL model. And then sample 20 shapes from them via FPS in the pose parameters of generation. Subsequently, similar to the procedure in Sec.4.2 of the main text, we construct 50 pairs among the 20 shapes for inference. As shown in Tab. 5 of the main text, our method is outperformed by SmoothShells with a noticeable margin.
To investigate the failure cause, we examine the intermediate results of our pipeline. As shown in Fig. 15, when there are significant differences in the pose, the image interpolation module (Zhang et al., 2023) struggles to return plausible results (top row), which is further amplified in the follow-up point cloud reconstruction (bottom row). Such discrepancy then leads to the suboptimal solution of our method on this benchmark.
@@ A.7. Running Time analysis
In particular, we evaluate MapTree  [^1] https://github.com/llorz/SGA20_mapExplor, SmoothShells  [^2] https://github.com/marvin-eisenberger/smooth-shells, NeuroMorph  [^3] https://github.com/facebookresearch/neuromorph, and BIM on all the test data with the regarding official implementations. Regarding ENIGMA, since the code is not publicly available, we asked the authors to run baseline evaluation, who also reported that ENIGMA took 20 mins to match shapes of 5000 vertices with RHM (Ezuz et al., 2019) as post-refinement.